Would you mind answering this question with clear writing and circle the correct number?.
The question contains a,b, and c,
A) please do it step by step.
B) so I can do my self thanks.

Your company is exploring investment opportunities for $15,000,000. The company would like to maximize the return on this investment. Four investment opportunities are being explored for a period of one year, as shown below. ii. iii. iv. 10% annual interest, compounded quarterly 9.85% annual interest, compounded daily 9.95% annual interest, compounded monthly 9.80% annual interest, compounded annually a.) Which opportunity should your company choose? How much will this opportunity earn? b.) What is the nominal interest rate of the best opportunity? c.) What is the effective interest rate of the best opportunity?
Answer #1

a> When a company chooses an investment, they should choose the one with the best effective interest rate.

For i, the effective interest rate is (1+0.1/4)^4-1=10.38%

For ii, the effective interest rate is (1+0.0985/365)^365-1=10.35%

For iii, the effective interest rate is (1+0.0995/12)^12-1=10.42%

For iv, the effective interest rate is 0.098)=9.8%

So, the company should invest in option 3 and they will earn $15,000,000x1.1041654=$16,562,481 after a year.

b> The nominal interest for that is 9.95%

c> The effective interest rate is 10.41654%